The average atomic weight of copper is 63.5 amu. Copper has two naturally occurring isotopes. One of theisotopes has an atomic weight of 62.9 and an abundance of 69.1%. The other isotope has an abundance of30.9%. What is the atomic weight of this isotope?
1
Expert's answer
2020-10-14T05:46:13-0400
decision
Ar (average) = Ar(1)* W +Ar (2) * W
Ar (1) = 62,9  W = 69,1% = 0,691
Ar(2)= Х   W=30,9% = 0,309

63,5=62,9 * 0,691 + Х * 0,309

63,5= 43,4639 + 0,309Х

0,309Х = 20,0361

Х = 64,84


answer: 64,84